OpenStudy (anonymous):

How do I solve this 3(4)^(3/2) ?

OpenStudy (jdoe0001):

\(\Large \bf a^{\frac{{\color{blue} n}}{{\color{red} m}}} = \sqrt[{\color{red} m}]{a^{\color{blue} n}} \qquad \qquad \sqrt[{\color{red} m}]{a^{\color{blue} n}}=a^{\frac{{\color{blue} n}}{{\color{red} m}}}\)

OpenStudy (anonymous):

so